please help.
The difference of two numbers is 6. The second is 9 less than 2 times the first. What are the two numbers?

Answer by tutorcecilia(2152)

x=first number
y=second number
difference = to subtract
9 less than = subtract 9
2 times = multiply by 2
is = equals
.
x-y=6
x-(2x-9)=6 [becasue the second number(y) is 9 less than twice the first (2x)]
.
x-(2x-9)=6 [solve for the x-term]
x-2x+9=6
-x+9=6
-x-9=6-9
-x=-3
x=3
.
check by plugging (x=3) into the original equation and solve:
x-(2x-9)=6
(3)-(2(3)-9)=6
3-(6-9)=6
3-(-3)=6
6=6 [checks out]